OpenStudy (anonymous):

how would you find the intersection between y=cosx and y=sin2x..... thanks :)

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

myininaya (myininaya):

set them equal cosx=sin2x cosx=2sinxcosx 1=2sinx 1/2=sinx does this help?

OpenStudy (anonymous):

careful, you forgot about when cos(x)=0

OpenStudy (anonymous):

thanks, but i was just wondering.....how did you get 1 on the left side

OpenStudy (anonymous):

It is better to pull the 2sinxcosx to the left and factor

myininaya (myininaya):

when is sinx 1/2? when x=pi/6 and when x=5pi/6 so the answer keeps going so x=pi/6+2npi for n=0,1,2,3,... and x=5pi/6+2npi for n=0,1,2,3,... i divided both sides by cosx to get 1 on that one side

myininaya (myininaya):

he right to get all of our intersections we need to bring everything on one side cosx-2sinxcosx=0 cosx(1-2sinx)=0 cosx=0 1-2sinx=0 => 1/2=sinx (we already solved this one) cosx=0 when x=pi/2+2npi and x=3pi/2+2npi n=0,1,2,...

myininaya (myininaya):

nice call oser

OpenStudy (anonymous):

thank you so much oser and myininaya

myininaya (myininaya):

np

OpenStudy (anonymous):

here's the whole kit and kaboodle... (see attached)

OpenStudy (anonymous):

thank you so much, that was really nice of you :D

OpenStudy (anonymous):

no problem... have a great one! :o)

OpenStudy (anonymous):

oser....another question.....how come u multiplied 2sinx with cosx on the right side

OpenStudy (anonymous):

one of the trig identities you should know is that sin(2x) = 2sin(x)cos(x). This comes from the addition identity for sines: sin(a+b)=sin(a)cos(b)+sin(b)cos(a). When both a and b equal x, you get sin(2x) = sin(x)cos(x) + sin(x)cos(x) = 2sin(x)cos(x).

OpenStudy (anonymous):

oh, wow, i did not know that rule...no wonder i was so confused..thank you so much! u were really helpful

OpenStudy (anonymous):

my pleasure. :o)